# GA-SmaAt-GNet: Generative Adversarial Small Attention GNet for Extreme Precipitation Nowcasting

### v1

@@ Abstract
In recent years, data-driven modeling approaches have gained considerable traction in various meteorological applications, particularly in the realm of weather forecasting. However, these approaches often encounter challenges when dealing with extreme weather conditions. In light of this, we propose GA-SmaAt-GNet, a novel generative adversarial architecture that makes use of two methodologies aimed at enhancing the performance of deep learning models for extreme precipitation nowcasting. Firstly, it uses a novel SmaAt-GNet built upon the successful SmaAt-UNet architecture as generator. This network incorporates precipitation masks (binarized precipitation maps) as an additional data source, leveraging valuable information for improved predictions. Additionally, GA-SmaAt-GNet utilizes an attention-augmented discriminator inspired by the well-established Pix2Pix architecture. Furthermore, we assess the performance of GA-SmaAt-GNet using real-life precipitation dataset from the Netherlands. Our experimental results reveal a notable improvement in both overall performance and for extreme precipitation events. Furthermore, we conduct uncertainty analysis on the proposed GA-SmaAt-GNet model as well as on the precipitation dataset, providing additional insights into the predictive capabilities of the model. Finally, we offer further insights into the predictions of our proposed model using Grad-CAM. This visual explanation technique generates activation heatmaps, illustrating areas of the input that are more activated for various parts of the network.
@@ 1 Introduction
Precipitation nowcasting nowcasting , also referred to as short-term precipitation forecasting, has long been a focal point of meteorological research. Especially, the ability to accurately predict short-term extreme precipitation events holds particular significance, as these weather phenomena have a substantial impact on both human life and economy. Notably, the Netherlands has experienced a noteworthy increase in precipitation over the past century, as highlighted in the findings of studies such as precip_trends_nl ; kmni_extreme . Furthermore, the frequency of extreme precipitation events in the Netherlands has been on the rise, as demonstrated by extreme_precip_nl ; kmni_extreme . In light of these developments, precise nowcasts for extreme precipitation events are of great importance for timely warnings and precautionary measures. Traditionally, meteorologists have relied upon Numerical Weather Prediction (NWP) models for generating weather forecasts nwp . NWP models employ intricate mathematical equations to simulate various atmospheric properties. This involves solving a set of complex partial differential equations governing the physical state of the atmosphere. While NWP models can deliver accurate forecasts, their computational demands make them less suitable for nowcasting tasks which require short-term forecasts, typically ranging from minutes to a maximum of six hours. In recent years, Convolutional Neural Networks (CNNs), a class of deep learning models particularly well-suited for image and spatial data analysis, have emerged as a promising area of research for tackling the precipitation nowcasting problem. CNNs have demonstrated remarkable success in various fields, including computer vision alex ; resnet ; vgg ; deep_shared , natural language processing kim2014convolutional ; zhang2016characterlevel , and, increasingly, meteorology convLSTM ; trajGRU ; deep_shared . Their ability to extract spatial patterns and features from radar or satellite imagery has made them valuable tools in the quest for more accurate precipitation forecasts, even for extreme weather events dgmr . However, it is essential to acknowledge that CNNs, while showing promise, often encounter significant challenges when predicting extreme precipitation events extreme_hard . These events exhibit complex and non-linear behaviors that can strain the capabilities of traditional deep learning models. As such, ongoing research in this field is not only focused on leveraging CNNs but also on enhancing their performance in predicting extreme weather events through innovative architectures dgmr , data augmentation techniques gan_mask , and the integration of additional meteorological variables nowcastnet . By addressing these challenges, meteorologists aim to harness the full potential of deep learning techniques in advancing the accuracy and reliability for extreme precipitation nowcasting.
In this paper, we present GA-SmaAt-GNet, a novel generative adversarial architecture designed to enhance the performance of the widely adopted SmaAt-UNet architecture SmaAtUnet in the context of extreme precipitation nowcasting. Our approach leverages a novel SmaAt-GNet architecture as the generator and incorporates an attention-augmented discriminator to achieve improved performance. Additionally, we provide a look into uncertainty of the models as well as the used precipitation data. Furthermore, we utilize Grad-CAM, a visual explanation technique, to generate heatmaps of the activations of different parts of our proposed network to gain further insight into its predictions. It is important to note that the precise definition of extreme precipitation can vary between different regions and research studies. For the purposes of this paper, when we refer to extreme precipitation in the Netherlands, we are specifically referring to precipitation depths exceeding 20 mm accumulated within a one-hour timeframe, as most drainage systems in the Netherlands are not able to process more than this amount of rain per hour knmi_zwaar .
@@ 2 Related work
In recent years, deep learning techniques have found applications in various weather forecasting tasks deep_shared . For instance, Shi et al. introduced ConvLSTM, which expanded upon the LSTM architecture by incorporating convolutional structures in the input-to-state and state-to-state transitions convLSTM . Their approach demonstrated improved performance when compared to traditional optical flow-based methods for precipitation nowcasting. Building upon the success of ConvLSTM, a subsequent study introduced TrajGRU trajGRU . TrajGRU extended ConvLSTM by employing a GRU (Gated Recurrent Unit) instead of an LSTM, a modification that enables the model to learn location-specific patterns more effectively.
@@
More recently, U-Net-based architectures have gained traction in the field of weather forecasting. Originally designed for biomedical image segmentation unet , the U-Net architecture has been repurposed effectively for various applications, including weather forecasting. This architecture consists of an encoder network, which progressively reduces the spatial resolution of the input image, and a decoder network that upsamples the feature maps back to the original image dimensions. These two networks are interconnected via skip connections, enabling the decoder to access high-resolution features from the encoder. While initially intended for biomedical image segmentation, the U-Net architecture has found successful application in various domains, including weather forecasting aatransunet ; SmaAtUnet ; sar . Trebing et al. SmaAtUnet extended the U-Net architecture by incorporating attention modules and depthwise-separable convolutions, successfully deploying it for precipitation nowcasting in the Netherlands and cloud cover nowcasting in France. The prediction performance of their SmaAt-UNet architecture is on par with the traditional U-Net, but with a notable reduction in the number of trainable parameters.
@@
Han et al. conducted a comprehensive study, as documented in precip_models_comparison , to assess the performance of various models in the radar precipitation nowcasting task. Among the models they examined were ConvLSTM and U-Net architectures. Their evaluation encompassed a period spanning from 2018 to 2020, as well as a specific case study focusing on a summer heavy rainfall event. The results of their investigation revealed a consistent trend: U-Net-based architectures consistently outperformed other models in terms of the critical success index (CSI) in both the long-term period and the context of the summer heavy rainfall event. Furthermore, their findings indicated that U-Net architectures achieved their best performance with shorter input sequences (60 minutes) compared to longer sequences (120 minutes). Additionally, they highlighted the significant influence of the most recent input frame on the accuracy of predictions.
@@
Recent research has also demonstrated the potential for enhancing forecast accuracy by integrating different types of data sources. For instance, Ko et al. radar_and_ground illustrated that combining radar-based precipitation data with ground-based weather data yields improved forecast performance. Additionally, Kaparakis et al. introduced WF-UNet wf-unet , highlighting the benefits of incorporating wind radar data into deep learning models for more accurate precipitation forecasts. The approach by Kaparakis et al. involved a two-stream architecture, with each stream featuring a distinct U-Net. One stream takes precipitation maps as input, while the other processes wind speed maps. The outputs from both streams are combined, and a 1x1 3D convolution is applied, followed by linear activation to generate the final prediction. To evaluate the effectiveness of their approach, they compared it with four other models, including the persistence model, core UNet unet , AsymmetricInceptionRes3DDR-UNet AsymmInceptionRes-3DDR-UNet , and Broad-UNet broadUnet , which did not incorporate wind data. Their findings conclusively demonstrated that WF-UNet outperformed all other models, highlighting the significant improvement in prediction performance achieved through the inclusion of wind data.
@@
Generative adversarial models (GANs) also offer significant potential for enhancing precipitation forecasts, particularly for higher rain intensities and for generating clearer predictions. Ravuri et al. introduced a novel approach called the deep generative model of radar (DGMR) dgmr , which demonstrated improved performance in these aspects compared to the conventional UNet framework. DGMR leverages a generator that utilizes convolutional gated recurrent units (GRUs) for enhanced modeling. Additionally, DGMR uses two discriminators, each focusing on distinct aspects of image quality assessment. One discriminator evaluates the temporal consistency, ensuring that predictions maintain a coherent progression over time. The other discriminator examines spatial consistency, ensuring that the generated images align with spatial patterns accurately. This approach contributes to more accurate and less blurry precipitation forecasts, particularly for higher rain intensities.
@@
For any warning system, understanding the confidence level of predictions is crucial. Unlike Bayesian models, deep learning tools lack a solid mathematical foundation for assessing model uncertainty. Foldesi et al. uncertainty_comparison investigated this issue by evaluating various methods for quantifying uncertainty in deep learning models, particularly when dealing with time-series data. Their findings highlighted two effective approaches: ensemble methods uncertainty_ensemble and Dropoutdropout /DropConnectuncertainty_dropconnect techniques. Ensemble methods involve training multiple copies of the same model architecture and combining their predictions. Lakshminarayanan et al. uncertainty_dropconnect demonstrated that ensembles offer reliable uncertainty quantification properties, making them a valuable choice. On the other hand, the dropout approach generates a posterior distribution during testing by running the model for a set number of epochs ($k$) for each target precipitation map. An uncertainty map is then derived by calculating the variance across all $k$ precipitation maps. This method has been effectively employed in studies such as dropout_brain and aatransunet . DropConnect is comparable to Dropout, with the distinction that it sets weights to zero rather than activations. These methods provide valuable tools for assessing uncertainty in deep learning models, a critical aspect for the reliability of warning systems.
@@ 3 Methods
We propose the GA-SmaAt-GNet architecture, a generative adversarial model that builds upon the SmaAt-UNet architecture SmaAtUnet . The SmaAt-UNet architecture, an extension of the UNet model, enhances the original design by introducing Convolutional Block Attention Modules (CBAM) cbam in the encoder to emphasize important features and depthwise-separable convolutions to reduce parameter count. These enhancements make SmaAt-UNet particularly well-suited for time series prediction tasks, facilitating precise value predictions for each pixel over time. The model’s efficiency and attention-driven feature selection contribute to its improved performance. In what follows, we will first introduce our proposed generator architecture, SmaAt-GNet. Then we will explain the loss function and our proposed discriminator architecture.
@@ 3.1 Proposed GA-SmaAt-GNet
We first introduce SmaAt-GNet (see Fig. 0(b)), which will serve as the generator of GA-SmaAt-GNet shown in Fig. 1. SmaAt-GNet builds upon SmaAt-UNet by incorporating a secondary stream encoder. This encoder facilitates the integration of an extra data source, specifically binary precipitation masks (refer to Section 4.2 for a comprehensive explanation), into the model. This additional encoder has the same architecture as the original SmaAt-UNet encoder. The CBAM layer results at each network level of the second encoder stream are concatenated with the CBAM layer results from the original encoder at the same network levels. These concatenated results are then connected to the decoder through skip connections. This design ensures that each network level can leverage features from both data sources, while also enabling attention-based feature selection within each encoder. By sharing the same decoder, the network can uncover relationships between features from both streams, while the features themselves are learned independently. We also include two dropout layers with a dropout probability of 0.5 after the first two bilinear upsampling layers in the decoder to prevent overfitting and they also allow us to approximate the model’s uncertainty as described in Section 5.2.1.
@@
where $x_{i}$ represents the input sequence, $y_{i}$ is the ground truth, $m_{i}$ are the precipitation masks, $n$ is the number of samples, $[D(\cdot,\cdot)]_{j}$ indicates the $j$th element of the $4\times 4$ matrix $[D(\cdot,\cdot)]$. The generator ($G$) seeks to minimize $\mathcal{L}_{\text{cGAN}}$, while the discriminator ($D$) aims to maximize it. Previous methodologies have demonstrated the advantage of combining the GAN objective with a regularization term, such as L1 or L2 distance pix2pix ; context_enc . In our case, we opt for L2 regularization over L1 regularization, given its greater sensitivity for extreme cases due to its squaring operation, which aligns with our focus.
@@
where $\lambda$ is a hyperparameter used to balance the GAN loss and regularization term.
It is crucial to note that we apply the regularization term exclusively when optimizing the generator, while for the discriminator, the default $\mathcal{L}_{\text{cGAN}}$ from Eq. (1) is optimized.
The dropout layers in the SmaAt-GNet architecture prevent the generator from becoming a simple deterministic mapping function. An alternative method to introduce randomness is by concatenating an image with random noise in the model. However, as noted in pix2pix , a model may simply learn to ignore this noise.
Our proposed discriminator architecture is shown in Fig. 0(c) and is an attention-augmented version of the Pix2Pix’s PatchGAN discriminator pix2pix . The PatchGAN architecture consists of several convolutional layers that process the input image in a hierarchical manner. Instead of classifying the entire image as real or fake, PatchGAN focuses on classifying small, overlapping patches of the input image. It also uses convolutions with a stride of two to gradually reduce the spatial resolution of the feature maps, helping it to capture information at different scales. We extend the PatchGAN architecture by adding an additional convolution before each convolution in the default PatchGAN architecture to prevent underfitting of the discriminator caused by a lack of model complexity. These added convolutions have a stride of one so that the spatial resolution does not decrease to quickly. Furthermore, after each convolution, a Convolutional Block Attention Module (CBAM) layer is added to focus on relevant parts of the features generated by the preceding convolutional layer.

### v2

@@ Abstract
In recent years, data-driven modeling approaches have gained significant attention across various meteorological applications, particularly in weather forecasting. However, these methods often face challenges in handling extreme weather conditions. In response, we present the GA-SmaAt-GNet model, a novel generative adversarial framework for extreme precipitation nowcasting. This model features a unique SmaAt-GNet generator, an extension of the successful SmaAt-UNet architecture, capable of integrating precipitation masks (binarized precipitation maps) to enhance predictive accuracy. Additionally, GA-SmaAt-GNet incorporates an attention-augmented discriminator inspired by the Pix2Pix architecture. This innovative framework paves the way for generative precipitation nowcasting using multiple data sources. We evaluate the performance of SmaAt-GNet and GA-SmaAt-GNet using real-life precipitation data from the Netherlands, revealing notable improvements in overall performance and for extreme precipitation events compared to other models. Specifically, our proposed architecture demonstrates its main performance gain in summer and autumn, when precipitation intensity is typically at its peak. Furthermore, we conduct uncertainty analysis on the GA-SmaAt-GNet model and the precipitation dataset, providing insights into its predictive capabilities. Finally, we employ Grad-CAM to offer visual explanations of our model’s predictions, generating activation heatmaps that highlight areas of input activation throughout the network.
@@ 1 Introduction
Precipitation nowcasting [1], also referred to as short-term precipitation forecasting, has long been a focal point of meteorological research. Especially, the ability to accurately predict short-term extreme precipitation events holds particular significance, as these weather phenomena have a substantial impact on both human life and economy. Notably, the Netherlands has experienced a noteworthy increase in precipitation over the past century, as highlighted in the findings of studies such as [2, 3]. Furthermore, the frequency of extreme precipitation events in the Netherlands has been on the rise, as demonstrated by [4, 3]. In light of these developments, precise nowcasts for extreme precipitation events are of great importance for timely warnings and precautionary measures. Traditionally, meteorologists have relied upon Numerical Weather Prediction (NWP) models for generating weather forecasts [5]. NWP models employ intricate mathematical equations to simulate various atmospheric properties. This involves solving a set of complex partial differential equations governing the physical state of the atmosphere. While NWP models can deliver accurate forecasts, their computational demands and limitations in resolving small-scale complex atmospheric processes [6] make them less suitable for nowcasting tasks which require short-term forecasts, typically ranging from minutes to a maximum of two hours. In recent years, Convolutional Neural Networks (CNNs), a class of deep learning models particularly well-suited for image and spatial data analysis, have emerged as a promising area of research for tackling the precipitation nowcasting problem. CNNs have demonstrated remarkable success in various fields, including computer vision [7, 8, 9], natural language processing [10, 11], cyber security [12, 13], and, increasingly, meteorology [6, 14, 15, 16, 17, 18]. Their ability to extract spatial patterns and features from radar or satellite imagery has made them valuable tools in the quest for more accurate precipitation forecasts, even for extreme weather events [19]. However, it is essential to acknowledge that CNNs, while showing promise, often encounter significant challenges when predicting extreme precipitation events [20]. These events exhibit complex and non-linear behaviors that can strain the capabilities of traditional deep learning models. As such, ongoing research in this field is not only focused on leveraging CNNs but also on enhancing their performance in predicting extreme weather events through innovative architectures [19], data augmentation techniques [21], and the integration of additional meteorological variables [22]. By addressing these challenges, meteorologists aim to harness the full potential of deep learning techniques in advancing the accuracy and reliability for extreme precipitation nowcasting.
It is important to note that the precise definition of extreme precipitation can vary between different regions and research studies. For the purposes of this paper, when we refer to extreme precipitation in the Netherlands, we are specifically referring to precipitation depths exceeding 20 mm accumulated within a one-hour time frame, as most drainage systems in the Netherlands are not able to process more than this amount of rain per hour [23].
In this study, we demonstrate the effectiveness of data augmentation and a generative approach for extreme precipitation nowcasting. To this end, we present GA-SmaAt-GNet, a novel generative adversarial framework designed to enhance the performance of the widely adopted SmaAt-UNet architecture [24] in the context of extreme precipitation nowcasting. Here, we briefly provide the key contributions of our work as follows:
  1. 1.
    We develop SmaAt-GNet which extends SmaAt-UNet by incorporating precipitation masks (binarized precipitation maps) as an additional data source. This provides supplementary information to the network and allows the model to better understand where areas of specific precipitation intensity are located. This architecture also paves the way for the integration of more data sources to enhance the predictive performance of deep learning models for meteorological tasks.
  2. 2.
    We develop GA-SmaAt-GNet, a generative adversarial framework, which utilizes our proposed SmaAt-GNet as a generator. Additionally, it incorporates a novel discriminator architecture that utilizes the CBAM attention mechanism [25]. This approach allows the discriminator to focus on critical areas in the data, improving its ability to distinguish between real and generated precipitation patterns.
  3. 3.
    We delve into both epistemic uncertainty of the models and aleatoric uncertainty inherent in the precipitation data. Our analysis reveals that both forms of uncertainty escalate with higher precipitation intensities, underscoring the inherent complexities involved in the task of extreme precipitation nowcasting.
  4. 4.
    We utilize Grad-CAM, a visual explanation technique, to generate heatmaps of the activations of different parts of our proposed network to gain further insight into its predictions. With these generated heatmaps we show that the precipitation maps and precipitation masks that serve as input to our proposed SmaAt-GNet architecture contribute to the distinct areas of the model’s prediction.
  5. 5.
    We evaluate the performance of our proposed models on a real-world precipitation dataset from the Netherlands. Spanning 25 years, from 1998 to 2022, this dataset provides a comprehensive and realistic environment to assess the model’s capabilities in predicting extreme weather events.
  6. 6.
    We conduct a comparative analysis of our proposed models against other established architectures, such as SmaAt-UNet [24] and RainNet [26]. Our results reveal that both proposed SmaAt-GNet and GA-SmaAt-GNet exhibit enhanced overall performance when evaluated on the precipitation dataset as well as a notable improvement in nowcasting extreme precipitation events.
@@ 2 Related work
In recent years, deep learning techniques have found applications in various weather forecasting tasks [6]. For instance, Shi et al. introduced ConvLSTM, which expanded upon the LSTM architecture by incorporating convolutional structures in the input-to-state and state-to-state transitions [14]. Their approach demonstrated improved performance when compared to traditional optical flow-based methods for precipitation nowcasting. Building upon the success of ConvLSTM, a subsequent study introduced TrajGRU [15]. TrajGRU extended ConvLSTM by employing a GRU (Gated Recurrent Unit) instead of an LSTM, a modification that enables the model to learn location-specific patterns more effectively.
@@
More recently, U-Net-based architectures have gained traction in the field of weather forecasting. Originally designed for biomedical image segmentation [27], the U-Net architecture has been repurposed effectively for various applications, including weather forecasting. This architecture consists of an encoder network, which progressively reduces the spatial resolution of the input image, and a decoder network that upsamples the feature maps back to the original image dimensions. These two networks are interconnected via skip connections, enabling the decoder to access high-resolution features from the encoder. While initially intended for biomedical image segmentation, the U-Net architecture has found successful application in various domains, including weather forecasting [28, 24, 29]. Trebing et al. [24] extended the U-Net architecture by incorporating attention modules and depthwise-separable convolutions, successfully deploying it for precipitation nowcasting in the Netherlands and cloud cover nowcasting in France. The prediction performance of their SmaAt-UNet architecture is on par with the traditional U-Net, but with a notable reduction in the number of trainable parameters.
@@
Han et al. conducted a comprehensive study, as documented in [30], to assess the performance of various models in the radar precipitation nowcasting task. Among the models they examined were ConvLSTM and U-Net architectures. Their evaluation encompassed a period spanning from 2018 to 2020, as well as a specific case study focusing on a summer heavy rainfall event. The results of their investigation revealed a consistent trend: U-Net-based architectures consistently outperformed other models in terms of the critical success index (CSI) in both the long-term period and the context of the summer heavy rainfall event. Furthermore, their findings indicated that U-Net architectures achieved their best performance with shorter input sequences (60 minutes) compared to longer sequences (120 minutes). Additionally, they highlighted the significant influence of the most recent input frame on the accuracy of predictions.
@@
Recent research has also demonstrated the potential for enhancing forecast accuracy by integrating different types of data sources. For instance, Ko et al. [31] illustrated that combining radar-based precipitation data with ground-based weather data yields improved forecast performance. Additionally, Kaparakis et al. introduced WF-UNet [32], highlighting the benefits of incorporating wind radar data into deep learning models for more accurate precipitation forecasts. The approach by Kaparakis et al. involved a two-stream architecture, with each stream featuring a distinct U-Net. One stream takes precipitation maps as input, while the other processes wind speed maps. The outputs from both streams are combined, and a 1x1 3D convolution is applied, followed by linear activation to generate the final prediction. To evaluate the effectiveness of their approach, they compared it with four other models, including the persistence model, core UNet [27], AsymmetricInceptionRes3DDR-UNet [33], and Broad-UNet [34], which did not incorporate wind data. Their findings conclusively demonstrated that WF-UNet outperformed all other models, highlighting the significant improvement in prediction performance achieved through the inclusion of wind data.
Generative adversarial models (GANs) also offer significant potential for enhancing precipitation nowcast, particularly for higher rain intensities and for generating clearer predictions. Generative Adversarial Networks (GANs) are a powerful method in generative machine learning, introduced by Goodfellow et al. [35]. GANs consist of two networks: a generator and a discriminator. The generator creates synthetic data resembling real data, while the discriminator distinguishes between real and fake samples. They train together in a competitive manner, with the generator improving its output to fool the discriminator, and the discriminator enhancing its ability to detect fakes. This adversarial process drives the generator to produce more realistic samples.
@@
Ravuri et al. introduced a novel approach called the deep generative model of radar (DGMR) [19], which demonstrated improved performance for extreme precipitation nowcasting compared to the conventional UNet framework. DGMR leverages a generator that utilizes convolutional gated recurrent units (GRUs) for enhanced modeling. Additionally, DGMR uses two discriminators, each focusing on distinct aspects of image quality assessment. One discriminator evaluates the temporal consistency, ensuring that predictions maintain a coherent progression over time. The other discriminator examines spatial consistency, ensuring that the generated images align with spatial patterns accurately. This approach contributes to more accurate and less blurry precipitation forecasts, particularly for higher rain intensities.
@@
For any warning system, understanding the confidence level of predictions is crucial. Unlike Bayesian models, deep learning tools lack a solid mathematical foundation for assessing model uncertainty. Foldesi et al. [36] investigated this issue by evaluating various methods for quantifying uncertainty in deep learning models, particularly when dealing with time-series data. Their findings highlighted two effective approaches: ensemble methods [37] and Dropout[38]/DropConnect[39] techniques. Ensemble methods involve training multiple copies of the same model architecture and combining their predictions. Lakshminarayanan et al. [39] demonstrated that ensembles offer reliable uncertainty quantification properties, making them a valuable choice. On the other hand, the dropout approach generates a posterior distribution during testing by running the model for a set number of epochs ($k$) for each target precipitation map. An uncertainty map is then derived by calculating the variance across all $k$ precipitation maps. This method has been effectively employed in studies such as [40] and [28]. DropConnect is comparable to Dropout, with the distinction that it sets weights to zero rather than activations. These methods provide valuable tools for assessing uncertainty in deep learning models, a critical aspect for the reliability of warning systems.
@@ 3 Methods
We propose the GA-SmaAt-GNet architecture, a generative adversarial model that builds upon the SmaAt-UNet architecture [24]. The SmaAt-UNet architecture, an extension of the UNet model, enhances the original design by introducing Convolutional Block Attention Modules (CBAM) [25] in the encoder to emphasize important features and depthwise-separable convolutions to reduce parameter count. These enhancements make SmaAt-UNet particularly well-suited for time series prediction tasks, facilitating precise value predictions for each pixel over time. The model’s efficiency and attention-driven feature selection contribute to its improved performance. In what follows, we will first introduce our proposed generator architecture, SmaAt-GNet. Then we will explain the loss function and our proposed discriminator architecture.
@@ 3.1 Proposed GA-SmaAt-GNet
We first introduce SmaAt-GNet (see Fig. 1(b)), where “GNet” signifies a G-shaped architecture in contrast to the U-shaped design of SmaAt-UNet. SmaAt-GNet will serve as the generator of GA-SmaAt-GNet shown in Fig. 1. SmaAt-GNet builds upon SmaAt-UNet by incorporating a secondary encoder. This encoder facilitates the integration of an extra data source, specifically binary precipitation masks (refer to Section 4.2 for a comprehensive explanation), into the model. This additional encoder has the same architecture as the original SmaAt-UNet encoder, which has CBAMs as the final module at each level of the encoder. The outputs of the CBAMs of both encoder streams are concatenated at the same level of the network. These concatenated results are then connected to the decoder through skip connections. This design ensures that each network level can leverage features from both data sources, while also enabling attention-based feature selection within each encoder. By sharing the same decoder, the network can uncover relationships between features from both streams, while the features themselves are learned independently. We also include two dropout layers with a dropout probability of 0.5 after the first two bilinear upsampling layers in the decoder to prevent overfitting and they also allow us to approximate the model’s uncertainty as described in Section 5.2.1.
@@
where $x_{i}$ represents the input sequence, $y_{i}$ is the ground truth, $m_{i}$ are the precipitation masks, $n$ is the number of samples, $[D(\cdot,\cdot)]_{j}$ indicates the $j$th element of the $4\times 4$ matrix $[D(\cdot,\cdot)]$. The generator ($G$) seeks to minimize $\mathcal{L}_{\text{cGAN}}$, while the discriminator ($D$) aims to maximize it. Previous methodologies have demonstrated the advantage of combining the GAN objective with a regularization term, such as L1 or L2 distance [41, 42]. In our case, we opt for L2 regularization over L1 regularization, given its greater sensitivity for extreme cases due to its squaring operation, which aligns with our focus.
@@
where $\lambda$ is a hyperparameter used to balance the GAN loss and regularization term. It should be noted that we apply the regularization term exclusively when optimizing the generator, while for the discriminator, the default $\mathcal{L}_{\text{cGAN}}$ from Eq. (1) is optimized.
The dropout layers in the SmaAt-GNet architecture prevent the generator from becoming a simple deterministic mapping function. An alternative method to introduce randomness is by concatenating an image with random noise in the model. However, as noted in [41], a model may simply learn to ignore this noise.
Our proposed discriminator architecture is shown in Fig. 1(c) and is an attention-augmented version of the Pix2Pix PatchGAN discriminator [41]. The PatchGAN architecture consists of several convolutional layers that process the input image in a hierarchical manner. Furthermore, instead of classifying the entire image as real or fake, PatchGAN focuses on classifying small, overlapping patches of the input image. It also uses convolutions with a stride of two to gradually reduce the spatial resolution of the feature maps, helping it to capture information at different scales. To prevent underfitting of the discriminator caused by a lack of model complexity, we extend the PatchGAN architecture by increasing the number of convolutions. These added convolutions have a stride of one so that the spatial resolution does not decrease to quickly. Furthermore, after each convolution, a Convolutional Block Attention Module (CBAM) layer is added to focus on relevant parts of the features generated by the preceding convolutional layer.